### v1

# From Ambiguity to Verdict: A Semiotic-Grounded Multi-Perspective Agent for LLM Logical Reasoning

Abstract
Logical reasoning is a fundamental capability of large language models (LLMs). However, existing studies largely overlook the interplay between logical complexity and semantic complexity, resulting in methods that struggle to address challenging scenarios involving abstract propositions, ambiguous contexts, and conflicting stances, which are central to human reasoning. For this gap, we propose LogicAgent, a semiotic-square–guided framework designed to jointly address logical complexity and semantic complexity. LogicAgent explicitly performs multi-perspective deduction in first-order logic (FOL), while mitigating vacuous reasoning through existential import checks that incorporate a three-valued decision scheme (True, False, Uncertain) to handle boundary cases more faithfully. Furthermore, to overcome the semantic simplicity and low logical complexity of existing datasets, we introduce RepublicQA, a benchmark that reaches college-level difficulty (FKGL = 11.94) and exhibits substantially greater lexical and structural diversity than prior benchmarks. RepublicQA is grounded in philosophical concepts, featuring abstract propositions and systematically organized contrary and contradictory relations, making it the most semantically rich resource for evaluating logical reasoning. Experiments demonstrate that LogicAgent achieves state-of-the-art performance on RepublicQA, with a 6.25% average gain over strong baselines, and generalizes effectively to mainstream logical reasoning benchmarks including ProntoQA, ProofWriter, FOLIO, and ProverQA, achieving an additional 7.05% average gain. These results highlight the strong effectiveness of our semiotic-grounded multi-perspective reasoning in boosting LLMs’ logical performance.
1 Introduction
Logical reasoning Smith (2003) plays a central role in human cognition, enabling structured transitions from ambiguous inputs to definitive conclusions. In AI Cohen et al. (2020); Vaswani et al. (2017), it underpins tasks such as commonsense reasoning Wang et al. (2024), mathematical proof Wang et al. (2023); Eisner et al. (2024); Gao et al. (2023), and philosophical thinking Paul (2013). However, robust logical reasoning in natural language remains challenging due to (1) semantic complexity Tuggy (1993), where expressions admit multiple interpretations or surface forms, and (2) logical complexity Gibson (1998), which requires reasoning over contextual premises and semantic interactions Zhang et al. (2020); Ding et al. (2024).
Recent approaches such as Chain-of-Thought (CoT) prompting Wei et al. (2022), Tree-of-Thought Yao et al. (2023), and Least-to-Most Zhou et al. (2022) have improved step-by-step reasoning, while symbolic-enhanced frameworks Pan et al. (2023); Ryu et al. (2024); Yang et al. (2023); Qi et al. (2025) further combine large language models (LLMs) Achiam et al. (2023); Guo et al. (2025); Zhao et al. (2023b) with first-order logic (FOL) solvers. These methods achieve strong results on widely used benchmarks such as ProntoQA Saparov & He (2022), ProofWriter Tafjord et al. (2020), FOLIO Han et al. (2022) and ProverQA Qi et al. (2025). Nevertheless, existing benchmarks remain confined to relatively simple and determinate semantics, often centered on everyday scenarios with shallow relations or logic problems that lack intricate semantic structure. Consequently, current evaluations fail to capture the full complexity of real-world reasoning, where propositions frequently involve philosophical abstraction, opposing stances, and deeper latent logical relations.
Figure 1: Overview of LogicAgent and the proposed RepublicQA benchmark. (Top-left) RepublicQA features abstract, philosophical propositions from Plato’s Republic with diverse contextual premises, enabling multiple semantic interpretations. (Bottom-left) LogicAgent consists of three stages. (Top-right) A multi-step reasoning process explores contraries and contradictions when S1 is indeterminate. (Bottom-right) LogicAgent outperforms strong baselines across four benchmarks, demonstrating robust generalization in symbolic, multi-perspective reasoning.
To capture the complex semantics and deep logical relations, we draw inspiration from Greimas’ Semiotic Square Greimas et al. (1982); Greimas (1987; 1988), a structuralist framework that extends binary oppositions into a four-part structure. It encompasses both contraries (e.g., $S_{1}$ vs. $S_{2}$, which cannot both be true but may both be false under non-empty domains) and contradictions (e.g., $S_{1}$ vs. $\lnot S_{1}$, which cannot both be true or false). We migrate this semantic framework into classical FOL with additional constraints, enabling structured multi-perspective reasoning that captures both complex semantics and deep abstract logical relations.
Motivated by agent-based paradigms Hong et al. (2023), we propose LogicAgent, a semiotic-square-guided reasoning framework that automates multi-perspective deduction through a three-stage pipeline: (1) Semantic Structuring Stage constructs a semiotic square to generate perspective variants of a proposition, including its contradiction and contrary, laying the foundation for multi-perspective reasoning and reflection. (2) Logical Reasoning Stage formalizes the contextual premises and performs symbolic deduction along both the original and contradiction paths. (3) Reflective Verification Stage assesses the reasoning trajectory through logic-aware reflection and revises conclusions when inconsistencies arise. This design enables LogicAgent to emulate human-like reasoning while systematically addressing semantic ambiguity and logical complexity.
To address the limitations of existing datasets in evaluating both semantic and logical complexity, we introduce RepublicQA, a benchmark derived from classical philosophical concepts. RepublicQA is manually annotated and cross-validated by multiple annotators to ensure high quality and consistency. Unlike existing reasoning benchmarks such as ProofWriter Tafjord et al. (2020), ProntoQA Saparov & He (2022), FOLIO Han et al. (2022), and ProverQA Qi et al. (2025), which rely on simple templates and semantically straightforward content, RepublicQA is designed to capture semantic complexity through abstract philosophical concepts and logical complexity by reasoning over contradictory or opposing positions. It consistently exceeds existing benchmarks in reading difficulty, lexical diversity, and long-span expressiveness. These properties make RepublicQA a unique testbed for reasoning that requires sensitivity to linguistic subtleties and deep logical relationships.
Experimental results show that our method achieves state-of-the-art performance on RepublicQA, surpassing strong baselines across different backbone models with an average improvement of 6.25%. To further validate its generalization, we also evaluate LogicAgent on ProntoQA, ProofWriter, FOLIO and ProverQA, where it again achieves superior results with an average gain of 7.05%. These findings confirm the effectiveness of semiotic-grounded multi-perspective reasoning in enhancing LLMs’ logical capabilities under ambiguity and conceptual complexity.
2 Preliminaries
Reasoning under ambiguity often involves not only binary truth values but also conceptual oppositions such as contraries (just vs. unjust) and contradictions (true vs. false). Classical logical formalisms capture the latter but lack a systematic way to encode the former. To address this gap, we incorporate Greimas’ Semiotic Square as a bridging device: it provides a structured representation of semantic oppositions, which we then ground in FOL. This integration forms the basis of LogicAgent, allowing it to align natural language semantics with formal logical deduction.
Figure 2: Greimas’ Semiotic Square: illustrating contraries ($S_{1}$ vs. $S_{2}$), contradictions ($S_{1}$ vs. $\lnot S_{1}$, $S_{2}$ vs. $\lnot S_{2}$), and implications ($S_{1}$ $\Rightarrow$ $\lnot S_{2}$, $S_{2}$ $\Rightarrow$ $\lnot S_{1}$).
Greimas’ Semiotic Square. The Greimas’ Semiotic Square Greimas et al. (1982) is a foundational construct in structuralist semantics that organizes conceptual contraries and contradictions into a four-element structure, enabling fine-grained reasoning over meaning, opposition, and implication. In our work, we migrate this semantic structure into the setting of classical FOL, with additional constraints to ensure logical soundness. Specifically, we introduce an existential import check to avoid vacuous truth from material implication, and extend the evaluation space from binary {True, False} to a three-valued scheme {True, False, Uncertain}, which better reflects reasoning under ambiguity.
Let $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} denote a primary proposition. The structure of the semiotic square is illustrated in Figure 2:
  • $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$}: the contrary of $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$}. The relation ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\ \bot\ {\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}$ implies both cannot be true, but both may be false (subject to non-empty domain constraints).
  • $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$}: the contradictory of $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$}, satisfying the classical law of excluded middle: ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\leftrightarrow\lnot{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}$.
  • $\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}$}: the contradictory of $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$}.
Semantic Implication Theorem. If $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} are contraries, then within the semiotic square the following semantic implications hold:
${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\Rightarrow{\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}\quad\text{and}\quad{\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}\Rightarrow{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}.$ (1)
Proof. Assume ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{True}$. Since $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} are contraries, we obtain ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}=\text{False}$. By the definition of contradiction, ${\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}=\text{True}$. Thus, ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\Rightarrow{\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}$. Symmetrically, ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}\Rightarrow{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}$.
This structural implication mechanism allows the reasoning agent to verify the coherence of judgments made about a target proposition by leveraging the relational semantics encoded in the square.
Figure 3: Overview of the LogicAgent framework. The agent processes a natural language proposition through three modular reasoning stages. (1) Semantic Structuring Stage constructs a Greimas’ Semiotic Square, generating four interrelated propositions: the primary proposition $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}S_{1}$}, its contradiction $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\lnot S_{1}$}, the contrary $\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}S_{2}$}, and the contradiction of the contrary $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\lnot S_{2}$}. These are verified for FOL-consistency using a CFG-based parser. (2) Logical Reasoning Stage transforms the premises into FOL, plans deductive steps for each proposition, and performs symbolic reasoning to evaluate their answers. (3) Reflective Verification Stage adjudicates the final judgment via three procedures: Direct Resolution, applied when $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} offer a contradictory answer; Quick Reflection, used when either $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} or $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is uncertain; and Deep Reflection, used when both $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} yield the same value, requiring further validation through the semiotic implication relations involving $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} and $\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}$}.
3 Methodology
To emulate human-like logical reasoning from multiple perspectives, we propose LogicAgent, as shown in Figure 3. The framework consists of three core stages: (1) Semantic Structuring Stage, (2) Logical Reasoning Stage, and (3) Reflective Verification Stage. By integrating semiotic theory with classical logic under additional constraints, LogicAgent enables multi-perspective reasoning and reflection over conceptual structures. Each stage plays a distinct role in transforming linguistic input into structured reasoning: from semantic structuring, to symbolic deduction, and ultimately to reflective verification for consistency.
3.1 Task Definition
Given a set of natural language Premises $P=\{p_{1},p_{2},\dots,p_{n}\}$, where each $p_{i}$ denotes a logical statement, and a Proposition $Q$, the task is to determine the answer of $Q$ with respect to $P$, choosing one of three labels: True, False, or Uncertain.  [^1] ProntoQA is restricted to the classical two-valued setting (True/False), whereas RepublicQA, ProofWriter, FOLIO, and ProverQA additionally include Uncertain to handle indeterminate cases.
3.2 Semantic Structuring Stage
Given an input proposition $Q$, this stage first treats it as the primary proposition $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$}, preserving its original semantic stance. Based on $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$}, the stage jointly generates its contradictory $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$}, the corresponding contrary $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$}, and the contradiction of the contrary $\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}$}, each paired with a symbolic representation in FOL.
Contradictory Construction.
Given a natural-language proposition $S_{1}$, we first formalize it into a FOL expression. We then negate the entire formula to obtain $\lnot S_{1}$ and simplify it using standard equivalences (quantifier negation, De Morgan, implication, bi-implication), as summarized in Table 1 (column $S_{1}$ (simplified)”). Finally, the simplified form is mapped back into natural language, ensuring that $\lnot S_{1}$ is both syntactically valid in FOL and semantically a strict negation of $S_{1}$.
Contrary Construction.
We adopt the classical definition of contrariety: $S_{1}$ and $S_{2}$ cannot both be true but may both be false. To block vacuous truth, we enforce non-emptiness and relevance constraints; when unmet, the evaluation defaults to Uncertain. The construction follows three paths:
  • Strict contraries: For canonical templates that always yield valid contraries (Table 1, “Strict”), we apply symbolic rules directly.
  • Conditional contraries: For templates requiring constraints to avoid vacuity (Table 1, “Conditional”), we first generate candidate contraries $S_{2}$ (via rules/LLM). Each undergoes an existential import check against premises $P$ to ensure non-empty domains, satisfiable antecedents, and instantiable variables. Valid candidates are then tested for contrariety with $S_{1}$ and contextual relevance (e.g., accept $A\!\to\!\lnot B$ only if $\mathrm{Sat}_{P}(A)$ holds).
  • Beyond coverage: The six rules in Table 1 cover common logical forms. For other cases (nested or domain-specific), we rely on model-assisted generation with self-supervised validation.
Validation and Verification.
All candidate propositions ($S_{1}$, $\lnot S_{1}$, $S_{2}$, $\lnot S_{2}$) are validated through a three-stage pipeline, and only those passing all stages are retained for downstream reasoning:
  1. 1.
    Truth-table evaluation: FOL formulas are assigned truth values to check whether relations of contrariety ($S_{1}$ vs. $S_{2}$) and contradiction ($S_{1}$ vs. $\lnot S_{1}$, $S_{2}$ vs. $\lnot S_{2}$) are satisfied.
  2. 2.
    CFG-based validation: A context-free grammar (CFG) checker enforces syntactic correctness of all FOL expressions, guaranteeing well-formedness.
  3. 3.
    LLM verification: An LLM confirms semantic and structural consistency, ensuring that contraries and contradictories remain faithful to the intended meaning and relevant to the premises.
Table 1: Unified rules for constructing contraries and contradictories. $A,B$ denote arbitrary formulas (possibly nested); $\varphi(x)$ denotes a predicate with variable $x$. $\oplus$ is exclusive-or.
# $S_{1}$ $\lnot S_{1}$ $S_{2}$ Type Constraint
1 $\forall x\,\varphi(x)$ $\lnot\forall x\,\varphi(x)\ \equiv\ \exists x\,\lnot\varphi(x)$ $\forall x\,\lnot\varphi(x)$ Strict N/A
2 $A\wedge B$ $\lnot(A\wedge B)\ \equiv\ \lnot A\vee\lnot B$ $A\wedge\lnot B$ Strict N/A
3 $A\leftrightarrow B$ $\lnot(A\leftrightarrow B)\ \equiv\ A\oplus B$ $A\leftrightarrow\lnot B$ Strict N/A
4 $\exists x\,\varphi(x)$ $\lnot\exists x\,\varphi(x)\ \equiv\ \forall x\,\lnot\varphi(x)$ $\exists x\,\lnot\varphi(x)$ Conditional $D\neq\emptyset$ (non-empty domain)
5 $A\rightarrow B$ $\lnot(A\rightarrow B)\ \equiv\ A\wedge\lnot B$ $A\rightarrow\lnot B$ Conditional $\text{Sat}(A)$ (antecedent satisfiable)
6 $A\vee B$ $\lnot(A\vee B)\ \equiv\ \lnot A\wedge\lnot B$ $A\vee\lnot B$ Conditional $A=\mathbf{F}$ (antecedent exclusion)
3.3 Logical Reasoning Stage
This stage comprises three functional units: a Translator for premise formalization, a Planner for reasoning path construction, and a Solver for logical deduction.
Translator.
The translator converts natural-language premises into FOL. Instead of relying on open-ended prompting, this step is guided by a set of general mapping conventions that define how linguistic structures are aligned with logical forms. In particular:
  • Entities (objects, concepts) $\mapsto$ unary predicates, e.g., $Entity(x)$.
  • Actions or relations $\mapsto$ binary or $n$-ary predicates, e.g., $Action(a,x)$ or $Relation(y,x)$.
  • Roles or agents $\mapsto$ unary predicates over individuals, e.g., $Role(y)$.
  • Normative or evaluative properties (just, good, harmful) $\mapsto$ predicates over actions or states, e.g., $Just(a)$, $Good(x)$.
This mapping schema is benchmark-agnostic and applies uniformly across different benchmarks. Each translated formula is further validated by a CFG parser to ensure syntactic correctness, so even if predicate names differ across benchmarks, the logical structure remains well-formed.
Planner.
For a selected proposition from the semiotic square (e.g., $S_{1}$), the planner constructs a reasoning blueprint. It specifies the evaluation goal, selects relevant premises, and determines which reasoning rules (e.g., Modus Ponens, Modus Tollens, Conjunction, Generalization) could be applied. The planner may also outline potential counterexample checks and identify implicit contextual relations that, while not explicitly stated in the premises, are salient within the discourse background. Its output is a structured reasoning trajectory, but without issuing a verdict.
Solver.
The solver operationalizes the planner’s blueprint: it applies the designated reasoning rules to the given premises, performs deductions step by step, and generates intermediate conclusions. During this process, it verifies logical consistency and checks for contradictions or counterexamples. The solver outputs both a transparent reasoning trace and the final classification of the proposition as True, False, or Uncertain.
3.4 Reflective Verification Stage
This stage adjudicates the final judgment through a three-stage reflective process that ensures coherence among the answers of the semiotic square’s four propositions.
Direct Resolution.
When $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} produce complementary verdicts, such as ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{True}$ and ${\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}=\text{False}$, the stage directly adopts the answer of $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} as final. This scenario reflects a decisive and non-contradictory judgment grounded in the strict contradiction relationship between the proposition and its contradictory. The decision rule for this resolution strategy is defined as follows:
[…]
Quick Reflection.
When either $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} or its contradictory $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is labeled as Uncertain, the stage triggers quick reflection by forwarding the two verdicts and their reasoning traces into a large language model. The model analyzes the internal consistency of the deduction process and returns a refined judgment based on four reflection types:
Case 1: If ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\in\{\text{T, F}\},\;{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}=\text{U}$ Type 1: $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} correct $\Rightarrow\;$ Return ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}={\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}$ Type 2: $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} incorrect $\Rightarrow\;$ Return ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{U}$ Case 2: If ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{U},\;{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}\in\{\text{T, F}\}$ Type 3: $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} correct $\Rightarrow\;$ Return ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{U}$ Type 4: $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} correct $\Rightarrow\;$ Return ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\lnot({\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}})$
Deep Reflection.
When both $\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} and its contradictory $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} yield the same verdict (e.g., both True or both False), this creates a contradiction under standard logical assumptions. The stage enters Deep Reflection mode, leveraging the structured semantic relations provided by the semiotic square, in particular the implications ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\Rightarrow{\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}$ and ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}\Rightarrow{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}$, to adjudicate which prediction is more likely to be valid.
Case 1: Both ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{True},\;{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}=\text{True}$  → Solve $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} If ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}=\text{True}$: since ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}\Rightarrow{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}$$\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is correct → Return ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{False}$ Else → Solve $\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}$} If ${\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}=\text{False}$: since ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}\Rightarrow{\color[rgb]{0.8,0.4,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0.4,0}\lnot S_{2}}$$\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}$} is incorrect → Return ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{False}$ Else: Invoke Quick Reflection Case 2: Both ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{False},\;{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}=\text{False}$  → Solve $\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}$} If ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}=\text{True}$: since ${\color[rgb]{0,0.4,0.4}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4,0.4}S_{2}}\Rightarrow{\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}}$$\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\lnot S_{1}$} is incorrect → Return ${\color[rgb]{0,0,0.8}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.8}S_{1}}=\text{False}$ Else: Invoke Quick Reflection
4 RepublicQA Benchmark
Figure 4: Complexity metrics comparison. Red is our benchmark.
Current benchmarks remain largely template-based and semantically simple, offering limited coverage of abstraction and ambiguity. To fill this gap, we construct RepublicQA, a benchmark designed to capture philosophical complexity and multi-perspective reasoning beyond existing datasets.
Benchmark Construction. RepublicQA is built from Plato’s Republic, a foundational text rich in discussions of justice, morality, governance, and knowledge. Its dialogical structure naturally yields abstract propositions and counterarguments, well suited for reasoning tasks involving ambiguity and opposing stances. We extracted propositions and contextual premises manually, with double annotation by two graduate students to ensure logical and semantic consistency.
Complexity Comparison. Unlike benchmarks such as FOLIO, ProofWriter, and ProverQA that remain largely template-based and semantically simple, RepublicQA introduces abstract propositions with deeper logical dependencies, balanced True/False/Uncertain distributions, and contexts requiring the integration of multiple philosophical concepts. We measure complexity across five dimensions: syntactic difficulty (FKGL), vocabulary richness (TTR), lexical variety (MTLD), phrasal diversity (UBR), and vocabulary size. Figure 4 shows that RepublicQA leads on four dimensions, requiring college-level comprehension and exhibiting greater lexical and phrasal diversity, while ranking second in vocabulary size. These properties underscore its abstract nature and resistance to template fitting, making it a valuable benchmark for reasoning under ambiguity and abstraction. More details about the RepublicQA benchmark can be found in Appendix D.
5 Experiment
5.1 Settings
We evaluate our framework on two fronts. First, we assess its performance on our proposed RepublicQA benchmark using different baseline models, verifying that the benchmark is broadly applicable for benchmarking logical reasoning. Second, to test the generalizability of our method, we conduct evaluations on established logical QA benchmarks, including ProntoQA, ProofWriter, FOLIO, and ProverQA. The experimental setup includes the following components:
Benchmarks.
We evaluate on four established logical reasoning benchmarks: ProntoQA (Saparov & He, 2022) (5-hop subset), ProofWriter (Tafjord et al., 2020) (depth-5, OWA setting), FOLIO (Han et al., 2022) (full expert-curated split), and ProverQA (Qi et al., 2025) (hard split with 500 examples, 6–9 reasoning steps). Detailed benchmark descriptions are provided in Appendix C.1.
Baselines.
We compare LogicAgent with five representative baselines: Naive Prompting, Chain-of-Thought (CoT) (Wei et al., 2022), Logic-LM (Pan et al., 2023), SymbCoT (Xu et al., 2024b), and Aristotle (Xu et al., 2024a). Detailed baseline introductions are provided in Appendix C.2.
Model.
For RepublicQA, we evaluate with both the locally deployed qwen2.5:32b Yang et al. (2025) and GPT-4o Hurst et al. (2024), ensuring robustness across open and closed source LLMs. For other benchmarks (ProntoQA, ProofWriter, FOLIO, ProverQA), we adopt qwen2.5:32b as the base model. In all experiments, the decoding temperature is fixed at 0.
Symbolic Toolkit.
To verify the syntactic validity of logical forms, we employ the nltk Bird (2006) library for CFG-based structural checking during the FOL parsing stage.
Table 2: Performance comparison across RepublicQA and other logical reasoning benchmarks. Best results are in bold, second-best are underlined.
Method RepublicQA Other Benchmarks
Qwen2.5-32B $\uparrow$ GPT-4o $\uparrow$ Avg $\uparrow$ Pronto $\uparrow$ ProofWriter $\uparrow$ FOLIO $\uparrow$ ProverQA $\uparrow$ Avg $\uparrow$
Naive 68.50 74.00 71.25 82.00 59.17 60.29 39.60 60.27
CoT 72.00 75.00 73.50 92.40 63.17 68.42 47.20 67.80
Logic-LM 70.00 73.50 71.75 91.89 63.82 71.93 62.40 72.51
SymCoT 76.00 80.50 78.25 95.20 64.67 70.59 57.20 71.92
Aristotle 74.50 82.50 78.50 94.80 63.23 68.68 56.20 70.73
LogicAgent 82.50 87.00 84.75 97.80 71.95 79.90 68.60 79.56
(+6.50) (+4.50) (+6.25) (+2.60) (+7.28) (+7.97) (+6.20) (+7.05)
5.2 Comparison with SOTA
Table 2 presents the main results from which we can draw several observations.
Our RepublicQA highlights the unique challenges of semantic ambiguity. On RepublicQA, Logic-LM performs comparably to the naive baseline, indicating that tool-augmented methods bring little advantage when facing symbolic and semantic ambiguity. In contrast, our LogicAgent achieves the best performance on both Qwen2.5-32B (82.50) and GPT-4o (87.00), surpassing the strongest baseline by an average of 6.25 points. This confirms that RepublicQA effectively stresses reasoning under ambiguity, and that our method is best suited to address these challenges.
Our LogicAgent generalizes effectively across several mainstream reasoning benchmarks. LogicAgent achieves an average improvement of 7.05 points over the best baseline, with consistent gains on Pronto (+2.60), ProofWriter (+7.28), FOLIO (+7.97), and ProverQA (+6.20). These results show that the proposed framework transfers robustly beyond RepublicQA and delivers superior performance on diverse logical QA benchmarks.
Our LogicAgent is particularly effective in multi-hop settings with semantic ambiguity. Benchmarks such as FOLIO, ProverQA, and RepublicQA require long reasoning chains over context-sensitive propositions, which often hinder purely neural or template-based approaches. LogicAgent consistently achieves the best performance on these benchmarks, confirming its strength in handling complex reasoning under ambiguity.
5.3 Ablation Study
To evaluate the contribution of our method, we conduct three sets of ablation experiments aimed at answering the following key questions:
  1. 1.
    How effective is each stage in our reasoning framework? We ablate components to measure their contribution.
  2. 2.
    What is the impact of FOL representations and natural language descriptions on reasoning performance? We disable either modality to assess their relative importance.
  3. 3.
    How do different components affect computational efficiency? We compare per-sample runtime across ablation settings to study the trade-off between cost and accuracy.
Q1: Impact of Core Reasoning Stages.
To address Q1, we conduct ablation studies on each of the three core stages in our method. Specifically:
  • For Stage 1 (Semantic Structuring), we disable the construction of the Greimas semiotic square and retain only the proposition matching the original proposition.
  • For Stage 2 (Logical Reasoning), we remove the planning process and directly attempt to solve the proposition without intermediate step generation.
  • For Stage 3 (Reflective Verification), we remove both Quick Reflection and Deep Reflection, and instead apply a rule-based Direct Resolution mechanism. Specifically, the model selects the final verdict by combining the base resolution strategy 2 with the supplemental decision rule:
[…]
Table 3: Ablation results under different configurations.
Setting ProofW. FOLIO ProverQA RepublicQA* Avg
w/o Square 65.17 72.06 56.60 76.50 67.58
w/o Plan 62.17 69.61 75.00 72.00 69.70
w/o Reflect 67.50 76.12 63.40 78.50 71.38
Ours 71.95 79.90 68.60 82.50 75.74
The ablation results in Table 3 demonstrate that each component in our framework contributes meaningfully to the overall performance. Removing the semiotic square stage causes a substantial decline, with the average accuracy dropping from 75.74 to 67.58 (-8.16). This indicates that analyzing propositions from multiple semantic perspectives, including contraries and contradictions, is critical for handling complex meanings. Excluding the reflective verification stage results in a smaller decrease to 71.38 (-4.36), suggesting that earlier reasoning stages already yield relatively reliable conclusions. Interestingly, removing the planning stage improves performance on ProverQA (from 68.60 to 75.00, +6.40); however, this removal leads to sharp declines on the other three datasets (-9.78 on ProofWriter, -10.29 on FOLIO, -10.50 on RepublicQA), reducing the average (excluding ProverQA) from 77.78 to 70.94 (-6.84). We observed that ProverQA gold chains typically involve 6–9 reasoning steps, while our planner often generates trajectories exceeding 10 steps. This suggests the existence of a reasoning complexity threshold, beyond which over-extended reasoning depth may impair performance, pointing to an important direction for future research.
(a) Input modalities (FOL vs. Statement).
(b) Reasoning efficiency.
Figure 5: Ablation studies: (a) input modalities and (b) reasoning efficiency.
Q2: Impact of FOL and Natural Language Inputs.
To investigate the respective contributions of FOL and natural language inputs, we conduct an ablation study by removing either the FOL representations or the natural language statements from the model input. The results are visualized in Figure 5(a).
Across four reasoning benchmarks, removing FOL input causes substantial performance degradation. On ProofWriter and FOLIO, accuracy drops from 71.95 to 33.33 and from 79.90 to 34.31, showing that classical FOL representations are critical for symbolic reasoning. Even on ProntoQA, which mainly tests basic deduction, performance falls sharply from 97.80 to 49.00. By contrast, ProverQA shows only a modest decline, likely because its hardest split already involves long reasoning chains, where performance bottlenecks stem more from semantic complexity than from explicit FOL representation alone.
In contrast, RepublicQA shows a different pattern. When the natural language input is removed and only FOL is retained, the accuracy declines substantially from 82.50 to 55.50. Removing FOL leads to a smaller drop to 71.00. This suggests that RepublicQA relies more heavily on the rich semantic information embedded in natural language. As a multi-semantic benchmark grounded in philosophical discourse, it includes contextual nuances such as implicit assumptions, shifting definitions, and pragmatic dependencies, which are not fully captured by symbolic logic alone.
Nevertheless, the best performance on RepublicQA is still achieved when both modalities are used (82.50), confirming that FOL contributes essential symbolic constraints even in semantically complex scenarios. Taken together, these results demonstrate that natural language enhances interpretability and contextual understanding, while FOL ensures inferential precision. Their integration enables more accurate and robust reasoning across a wide range of tasks.
Q3: Impact on Computational Efficiency.
We measure per-sample processing time under various ablation settings on RepublicQA. As shown in Figure 5(b), removing planning achieves the largest speedup (-37.1%), reflecting the cost of orchestrating multi-step reasoning. Excluding FOL (woFOL) also reduces processing time by 35.4%, highlighting the overhead of symbolic deduction. Smaller reductions are observed when removing natural language statements (-15.5%) or reflective verification (-9.5%). Similar trends hold for ProofWriter, FOLIO and ProverQA, with planning and FOL as the main bottlenecks. For ProverQA, runtime is instead dominated by reasoning chain length, as removing FOL or statements has little effect.
6 Conclusion
We present LogicAgent, a reasoning framework designed to address linguistic ambiguity and complex semantic dependencies through structured, multi-perspective reasoning. Grounded in semiotic principles, LogicAgent transforms natural language into first-order logic representations and performs reasoning in three stages: semantic structuring of contraries and contradictions, FOL deduction along multiple reasoning paths, and reflective verification to refine conclusions. To evaluate reasoning under ambiguity and conceptual variability, we introduce RepublicQA, a benchmark derived from philosophical discourse with context-sensitive, semantically rich propositions. Compared with existing benchmarks such as ProntoQA, ProofWriter, FOLIO, and ProverQA, RepublicQA uniquely emphasizes symbolic–semantic alignment under ambiguity, where tool-augmented baselines show limited advantage. Experiments demonstrate that LogicAgent achieves the best performance on RepublicQA and also generalizes strongly across other benchmarks. These findings confirm both the distinct value of RepublicQA and the effectiveness of semiotic-grounded, logic-aware reasoning.
Ethics Statement
The RepublicQA benchmark introduced in this paper is derived entirely from publicly available philosophical texts, specifically Plato’s Republic, which is in the public domain. No human subjects, private data, or sensitive groups are involved, and the dataset poses no immediate ethical or societal risks. All processing steps, including proposition extraction and annotation, were carried out by graduate students following academic standards. For transparency and reproducibility, we plan to release both the code for dataset construction and the RepublicQA benchmark used in our experiments.
Reproducibility Statement
We provide partial code and benchmark with the submission. The complete implementation and the full RepublicQA benchmark will be released at the camera-ready stage to ensure full reproducibility. All experimental settings, model configurations, and benchmark processing details are documented in the main text and Appendix.
Appendix
This appendix provides supplementary materials, including the related work, formalization in first-order logic (FOL), details of baselines and benchmarks, the construction of RepublicQA, error analysis, case studies, computational efficiency evaluation, and full prompting examples.
Appendix A Related Work
LLM-Based Logical Reasoning. CoT prompting Wei et al. (2022); Zhang et al. (2025a); Mirzadeh et al. (2024) improves LLM reasoning Gu & Dao (2023); Bai et al. (2023); Huang & Chang (2022) by generating intermediate steps in natural language. Variants such as self-consistency Wang et al. (2022) and symbolic CoT Xu et al. (2024b) improve accuracy and interpretability, with SymbCoT showing that symbolic forms enhance logical reasoning. Aristotle Xu et al. (2024a) further introduces a logic-guided framework using decomposition, search, and resolution, achieving strong results on complex tasks. However, most CoT-based He et al. (2025) and symbolic methods follow linear reasoning paths Zheng et al. (2024); Zhao et al. (2023a) and struggle to capture semantic depth, including contraries and contradictions. Other approaches translate LLM outputs into external logic engines Pan et al. (2023); Ryu et al. (2024), but suffer from brittleness and lack of feedback. In contrast, we propose a fully internal symbolic framework in which the LLM constructs, manipulates, and verifies logical structures. Guided by Greimas’ semiotic square Greimas (1988), our method enables reasoning over semantically diverse, multi-perspective statements.
LLM-Powered Agents. Advances in LLMsSong et al. (2022; 2023); Li et al. (2024); Song et al. (2024); Hu et al. (2025); Song et al. (2025); Ye et al. (2025) have led to agent frameworks Zhang et al. (2025b); Durante et al. (2024) capable of planning, memory, and multi-step reasoning. Systems such as Generative Agents Park et al. (2023), AutoAgents Chen et al. (2023), and MetaGPT Hong et al. (2023) simulate interactive behavior or coordinate task execution, while others like Code-as-Policies Liang et al. (2023), Gorilla Patil et al. (2024), and TaskMatrix Liang et al. (2024) integrate APIs or GUI actions for real-world applications. Building on the agent paradigm, our approach leverages structured and automated reasoning with symbolic representation and reflective verification to tackle abstract and semantically diverse reasoning tasks. By centering the reasoning process on symbolic representation and multi-perspective analysis, we aim to extend the capabilities of LLMs without additional fine-tuning.
Benchmarks for Logical Reasoning. Existing logical reasoning benchmarks Patel et al. (2024) primarily focus on formal validity under controlled conditions. PrOntoQA Saparov & He (2022) is a synthetic benchmark targeting relational reasoning, often centered on transitivity and set membership in fully controlled symbolic settings. ProofWriter Tafjord et al. (2020) features synthetic natural language problems built around symbolic rulebases, where contexts typically involve simplified everyday entities such as animals, colors, and objects, and reasoning proceeds through connectives such as and, or, and not. FOLIO Han et al. (2022) introduces natural language problems annotated with FOL formulas, covering commonsense-style scenarios such as daily events and generic facts. ProverQA Qi et al. (2025) is built with the ProverGen framework, combining LLM generation with automated theorem proving to yield high-quality FOL reasoning tasks. It contains 1,500 dev examples across easy, medium, and hard splits (500 each), with the hard split emphasizing longer reasoning chains and formally verified conclusions. Despite their formal rigor, these benchmarks largely consist of concrete, unambiguous propositions with fixed semantic interpretations. They lack abstract concepts, contextual variability, and structured semantic contrasts such as contraries and contradictions, offering limited support for evaluating multi-perspective, contrastive reasoning over semantically rich and philosophically grounded domains.
Table 4: Key Syntax Elements in First-Order Logic
Name FOL Notation Explanation
Variable x,y,zx,\ y,\ z Placeholder symbols representing arbitrary elements in the domain of discourse.
Constant a,b,ca,\ b,\ c Refer to specific, fixed objects in the domain.
Operators (OP) $\{\oplus$, $\vee$, $\wedge$, $\rightarrow$, $\leftrightarrow$}\} Defines the set of logical connectives used to combine or relate propositions, including exclusive or, or, and, implication, and biconditional. Used in building compound formulas.
Function f(x),g(x,y)f(x),\ g(x,y) Maps input objects to an output object; returns a term.
Predicate P(x),R(x,y)P(x),\ R(x,y) Express properties or relations; returns true or false.
Negation $\lnot P(x)$ Logical NOT: $P(x)$ is not true.
Conjunction $P(x)\land Q(x)$ Logical AND: both $P(x)$ and $Q(x)$ must be true.
Disjunction $P(x)\lor Q(x)$ Logical OR: at least one of $P(x)$ or $Q(x)$ must be true.
Implication $P(x)\rightarrow Q(x)$ Logical implication: if $P(x)$ is true, then $Q(x)$ must be true.
Biconditional $P(x)\leftrightarrow Q(x)$ Logical equivalence: $P(x)$ and $Q(x)$ are true or false together.
Universal Quantifier $\forall x\;P(x)$ “For all $x$, $P(x)$ is true” — generalization.
Existential Quantifier $\exists x\;P(x)$ “There exists $x$ such that $P(x)$ is true” — existential claim.
Term $x$, $a$, $f(a,x)$ The basic expressions referring to objects (variables, constants, or functions).
Atomic Formula $P(a,x)$ A predicate applied to terms — indivisible logical unit.
Complex Formula $\forall x(P(x)\rightarrow Q(f(x)))$ A formula built from atoms using connectives and quantifiers.
WFF (Well-formed) A syntactically valid FOL formula interpretable as true or false.
Appendix B First-order Logic (FOL)
First-Order Logic (FOL), also known as predicate logic or first-order predicate calculus, is a formal system widely used in mathematics, computer science, philosophy, and linguistics. It extends propositional logic by introducing variables that range over objects in a domain and predicates that describe relationships and properties of these objects. FOL allows us to write general statements involving quantifiers, such as “for all” and “there exists,” making it a powerful tool for expressing logical structure and reasoning.
B.1 Formal Syntax and Validation of FOL
FOL forms the backbone of our symbolic reasoning pipeline. As shown in Table 4, FOL comprises several syntactic components that define the structure of logical statements, including variables, constants, predicates, logical operators, quantifiers, and term compositions.
FOL CFG Grammar. To ensure the well-formedness of FOL expressions, we implement a symbolic parser using the nltk library Bird (2006). Specifically, we define a context-free grammar (CFG) to support automatic parsing and validation of logical formulas throughout our pipeline:
S $\rightarrow$ F | Q F
Q $\rightarrow$ QUANT VAR | QUANT VAR Q
F $\rightarrow$$\lnot$’ ‘(’ F ‘)’ | ‘(’ F ‘)’ | F OP F | L
OP $\rightarrow\oplus$’ | $\vee$’ | $\wedge$’ | $\rightarrow$’ | $\leftrightarrow$
L $\rightarrow$$\lnot$’ PRED ‘(’ TERMS ‘)’ | PRED ‘(’ TERMS ‘)’
TERMS $\rightarrow$ TERM | TERM ‘,’ TERMS
TERM $\rightarrow$ CONST | VAR
QUANT $\rightarrow\forall$’ | $\exists$
Example: For the rule $\forall x(\text{Debt}(x)\wedge\text{Repaid}(x)\rightarrow\lnot\text{Just}(x))$”, the CFG derivation proceeds as follows, as shown in Figure 6:
  • QUANT $\rightarrow\forall$
  • PRED $\rightarrow$ ‘Debt’ | ‘Repaid’ | ‘Just‘
  • VAR $\rightarrow$ ‘x‘
Note that PRED, CONST, and VAR are instantiated dynamically for each example during parsing. This grammar enables symbolic structure checking and forms the foundation for all logic-based components in our agent.
Syntactic Validation. We incorporate a rigorous syntactic validation mechanism based on this CFG, serving as a critical quality control step prior to symbolic reasoning. The validator performs structural analysis to ensure:
  • Quantifier Scope Verification: Ensuring proper binding and scope relationships for universal and existential quantifiers
  • Predicate Structure Validation: Confirming syntactic correctness of predicate-argument structures
  • Logical Connective Placement: Verifying appropriate positioning and precedence of logical operators
Only expressions that pass CFG validation are forwarded to the reasoning phase. This ensures the logical integrity of FOL representations derived from natural language and prevents errors caused by malformed logical forms.
Figure 6: An example CFG parse tree for the FOL rule $\forall x(\text{Debt}(x)\wedge\text{Repaid}(x)\rightarrow\lnot\text{Just}(x))$.
Appendix C Benchmarks and Baselines
C.1 Benchmarks
ProntoQA is a synthetic question-answering benchmark designed to systematically explore the reasoning abilities of language models through formal analysis. The benchmark generates examples with chains-of-thought that describe the reasoning required to answer questions correctly, enabling systematic exploration of LLM reasoning capabilities. The benchmark focuses on fundamental logical relationships and deductive reasoning patterns, providing a controlled environment for assessing model performance on multi-step logical reasoning tasks.
ProofWriter is a synthetic benchmark featuring natural language problems that assess systematic neural logical deduction. Developed by the Allen Institute, ProofWriter generates implications, proofs, and natural language reasoning over rulebases of facts and rules under open world assumptions. This benchmark presents complex logical relationships involving combinations of conjunctions and disjunctions, requiring models to perform multi-step deductive reasoning while generating natural language proofs that justify their conclusions. And the context in this benchmark contains more challenging logical relationships such as the combination of ”and” and ”or.”
FOLIO is a natural language reasoning benchmark with fol reasoning problems that require models to determine the correctness of conclusions given a world defined by premises. FOLIO aims to ensure high language naturalness and complexity, an abundant vocabulary, and factuality while maintaining high reasoning complexity. It is a high-quality and manually curated benchmark, written by CS undergraduate and graduate students and researchers in academia and industry. To ensure that the conclusions follow the premises logically, all reasoning examples are annotated with FOL formulas. FOLIO represents one of the most challenging logical reasoning benchmarks, combining natural language complexity with the precision of FOL.
ProverQA is a high-quality FOL reasoning benchmark created with the ProverGen framework, which combines the generative diversity of LLMs with the rigor of automated theorem proving. Each instance includes natural language statements, FOL translations, and formally verified reasoning chains. The benchmark is designed to test deductive consistency and the ability to align symbolic and linguistic representations. The dev set contains 1,500 examples evenly divided into easy (1–2 reasoning steps), medium (3–5 steps), and hard (6–9 steps) levels, providing a scalable and systematically validated environment for evaluating logical reasoning under increasing complexity.
RepublicQA is a philosophical reasoning benchmark derived from Plato’s Republic, featuring complex logical reasoning problems that require models to evaluate the correctness of philosophical statements given contextual premises. RepublicQA aims to assess models’ capabilities in handling abstract philosophical concepts, moral reasoning, and classical logical arguments while maintaining high complexity in both language and reasoning patterns. The benchmark is carefully curated to reflect fundamental questions in western philosophy. Each example requires models to engage with sophisticated philosophical arguments, often involving conditional reasoning, moral principles, and abstract conceptual relationships. To ensure logical consistency, examples are structured around classical philosophical dialogues and arguments, requiring multi-step reasoning to determine whether conclusions follow from the given premises about justice, virtue, and political philosophy.
Figure 7: Answer distribution across different benchmarks.
C.2 Baselines
Here we illustrate the details of each baseline used for comparison.
Naive Prompting.
This approach involves directly presenting the question to the model without any additional instructions or intermediate reasoning steps. The model is expected to produce the final answer based solely on its pretrained knowledge. It does not encourage or structure the reasoning process, making it suitable only for simple or factual queries.
Chain-of-Thought (CoT).
Chain-of-Thought prompting introduces intermediate reasoning steps by encouraging the model to articulate its thought process before reaching a conclusion. Instead of predicting the final answer directly, the prompt guides the model to perform step-by-step reasoning. This has been shown to improve multi-step reasoning performance on complex tasks (Wei et al., 2022).
Logic-LM.
Logic-LM converts natural language statements into symbolic logic representations and uses a rule-based logical engine to carry out reasoning. By explicitly leveraging classical FOL rules, it enhances reasoning accuracy and structure, especially in tasks requiring strict logical consistency (Pan et al., 2023).
SymbCoT.
SymbCoT extends CoT prompting by incorporating symbolic representations and logic-based rules into the reasoning framework. This innovative approach first transforms natural language input into structured symbolic expressions, creating a FOL foundation for subsequent reasoning steps. The system then prompts the language model to reason over these symbolic representations with the assistance of FOL constraints and rules. This integration fundamentally improves the model’s capability for structured and formally grounded reasoning by bridging the gap between natural language understanding and symbolic logic (Xu et al., 2024b).
Aristotle.
Aristotle is a logic-complete reasoning framework designed for complex logical tasks by integrating symbolic structures across all reasoning stages. It includes three main components: Logical Decomposer, Logical Search Router, and Logical Resolver. These components enable structured task decomposition, guided reasoning, and contradiction resolution. Aristotle is particularly effective when used with LLMs, which benefit from symbolic supervision and structured guidance, leading to stronger performance than smaller models or purely generative strategies (Xu et al., 2024a).
Appendix D Details of Our RepublicQA
D.1 Statistics
Answer Distribution. The benchmark exhibits a balanced distribution across three answer categories, as illustrated in Figure 7. The relatively high proportion of ”Uncertain” answers (37.5%) reflects the nuanced nature of philosophical reasoning, where definitive conclusions are often difficult to establish.
Table 5: Basic Statistics of RepublicQA Benchmark
Metric Value
Total samples 200
Unique topics 58
Avg. context length 637.42
Avg. proposition length 166.42
Avg. explanation steps 6.10
Explanation completeness 100.0%
Basic Statistics. The RepublicQA benchmark comprises 200 carefully constructed samples covering 58 unique philosophical topics. Table 5 presents the fundamental statistical characteristics of the benchmark.
D.2 Philosophical Concepts
The RepublicQA benchmark is rich in philosophical content, drawing directly from the core themes of Plato’s Republic. An analysis of the frequency and interconnections of key philosophical concepts provides insight into the benchmark’s thematic depth. Figure 8(a) visualizes the frequency of these concepts.
As shown in the Figure 8(a), Justice is the most prevalent concept, appearing 1308 times, revealing its central role in the Republic. Other frequently occurring concepts include State (846 occurrences), Soul (670 occurrences), Art (451 occurrences), Knowledge (242 occurrences), Virtue (206 occurrences), and Education (117 occurrences). The frequent occurences of Justice and State reflect Plato’s exploration of the just city and the just individual.
D.3 Topic Modeling Results
We applied Latent Dirichlet Allocation (LDA) topic modeling to identify thematic structures within the RepublicQA benchmark. Using optimal hyperparameters determined through coherence score validation, we extracted five distinct thematic clusters that capture the core philosophical themes of Plato’s Republic:
  1. 1.
    Political Philosophy: Encompasses discussions of governance structures, including concepts such as tyrants, wealth distribution, rulers’ responsibilities, freedom, and systemic injustice.
  2. 2.
    Individual Psychology: Focuses on human nature and development, explaining personal desires, educational processes, external influences on character formation, and individual moral development.
  3. 3.
    Metaphysics and Epistemology: Examines questions about knowledge and reality, including the three parts of the soul, how we think rationally, and how different mental faculties work together.
  4. 4.
    Philosophy of Art and Reality: Addresses questions of representation and truth, examining concepts of essence, philosophical debate methodology, imitation theory, and the distinction between appearance and reality.
  5. 5.
    Ethics and Justice Theory: Examines moral frameworks, investigating justice principles, responses to injustice, ethical rules, power dynamics, and competing moral interests.
These thematic clusters demonstrate the benchmark’s comprehensive coverage of Platonic philosophy while maintaining balanced representation across major philosophical domains.
(a) Philosophical Concept Frequency
(b) Overall Word Cloud and Key Words
Figure 8: Analysis of Philosophical Concepts: (a) frequency distribution of concepts, (b) overall word cloud highlighting key terms.
D.4 Complexity Analysis of RepublicQA
RepublicQA presents a uniquely challenging benchmark for reasoning models due to its rich semantic diversity and structural complexity. Topic modeling reveals five core clusters corresponding to key themes in Platonic philosophy, ensuring evaluation across varied reasoning contexts rather than a single domain. This semantic distribution requires models to integrate multiple philosophical concepts simultaneously, making superficial pattern matching insufficient. In addition, the benchmark exhibits a balanced distribution of answer types and notable linguistic complexity, fostering evaluation scenarios that demand abstraction, disambiguation, and compositional reasoning.
We assess the linguistic and semantic complexity of RepublicQA by comparing it with four standard reasoning benchmarks: ProntoQA, FOLIO, ProofWriter, and ProverQA. Our analysis covers both surface and structural attributes using established metrics.
Linguistic Complexity. We compute the Flesch-Kincaid Grade Level (FKGL) to measure syntactic difficulty:
$\text{FKGL}=0.39\cdot\frac{N_{\text{words}}}{N_{\text{sentences}}}+11.8\cdot\frac{N_{\text{syllables}}}{N_{\text{words}}}-15.59$
Higher FKGL scores indicate more complex sentence structures and denser vocabulary.
Semantic Diversity. We report Type-Token Ratio (TTR) for vocabulary richness, MTLD for sustained lexical variety, and Unique Bigrams Ratio (UBR) for phrasal diversity. These metrics capture both abstraction and expression variability.
[…]
These results establish RepublicQA as a valuable resource for evaluating deep reasoning and generalization in artificial intelligence systems. Figure 8(b) further illustrates its conceptual landscape through a word cloud of prominent philosophical terms, reinforcing its role as a benchmark for semantically diverse and abstract reasoning tasks.
Table 6: Complexity Metrics Comparison (Higher is Better)
Benchmark FKGL $\uparrow$ TTR $\uparrow$ MTLD $\uparrow$ UBR $\uparrow$ Vocab
ProntoQA (Saparov & He, 2022) 6.78 0.448 13.93 0.852 69
ProofWriter (Tafjord et al., 2020) 1.25 0.193 11.31 0.513 61
FOLIO (Han et al., 2022) 6.62 0.569 33.54 0.805 1,021
ProverQA (Qi et al., 2025) 8.41 0.362 49.43 0.639 4,891
RepublicQA (Ours) 11.94 0.685 74.81 0.929 2,083
Appendix E Case Study
We present a representative case from RepublicQA to illustrate how LogicAgent integrates dual-form representations, multi-perspective reasoning, and reflective verification in a unified pipeline. We analyze this example through three key components of our methodology.
E.1 Dual-Form Representation and Semantic Precision
LogicAgent processes each proposition using both natural language and FOL representations. This dual-form design retains the conceptual richness of natural language while enabling symbolic reasoning under FOL. In the selected case, natural language captures nuanced distinctions (e.g., “justice” vs. “ability”), while FOL clarifies logical scope and reasoning structure. This integration supports:
  • Semantic Preservation: Contextual meaning is preserved during FOL translation
  • Logical Precision: Symbolic structure enables explicit reasoning
  • Boundary Clarification: FOL delineates abstract concepts
E.2 Multi-Perspective Reasoning for Robust Evaluation
To go beyond single-path deduction, LogicAgent constructs a semiotic square for each proposition, enabling reasoning over four semantic positions: $S_{1}$, $S_{2}$, $\lnot S_{1}$, and $\lnot S_{2}$. In this case, the system reasons over $S_{1}$ (“The just man is a thief”) and its contradiction $\lnot S_{1}$, revealing a conflict between their conclusions. This multi-perspective reasoning allows:
  • Verification Through Redundancy: Independent chains confirm or challenge conclusions
  • Error Detection: Logical inconsistency between perspectives triggers correction
  • Semantic Exploration: Opposing positions clarify conceptual boundaries
E.3 Planning, Execution, and Reflective Correction
Step 1 – Planning: A 7-step reasoning plan is generated for $S_{1}$ via semantic decomposition and rule mapping.
Step 2 – Reasoning Execution: $S_{1}$ yields an Uncertain result, while $\lnot S_{1}$ concludes True, signaling inconsistency.
[…]
Final Conclusion: The system resolves the inconsistency and outputs False for the original proposition “The just man turns out to be a thief”.
This case demonstrates how LogicAgent’s integrated reasoning pipeline, which combines dual-form representation, multi-perspective structure, and reflective correction, enables robust and explainable reasoning in semantically rich domains.
Appendix F Error Analysis
Figure 9: Overall and relation-specific accuracy across datasets.
Figure 10: Distribution of contradictory vs. contrary cases across datasets.
Our error analysis highlights four key capabilities required for strong logical reasoning: (1) accurate construction of Greimas’ semantic squares, (2) faithful FOL translation, (3) effective planning of reasoning paths, and (4) consistent verification.
Benchmark Semantic Richness Impact. As shown in Figure 10, the proportion of contrary relations varies widely across benchmarks. RepublicQA contains the richest semantic oppositions, with about 73% of instances involving meaningful contraries (e.g., justice vs. injustice). In contrast, FOLIO and ProverQA fall below 50%, reflecting limited availability of contraries despite their complexity. ProntoQA and ProofWriter contain almost none, restricting reasoning largely to contradictory checks. Notably, even in RepublicQA, certain propositions cannot yield valid contraries, underscoring that constructing oppositions is inherently difficult due to semantic ambiguity and hidden contextual links. Moreover, Figure 10 shows that accuracy on contrary cases is consistently lower than the overall accuracy, highlighting that contrary reasoning is intrinsically harder and more error-prone.
FOL Translation Accuracy. We observe relatively few FOL parsing errors with Qwen2.5-32B, especially on semantically rich datasets. However, even small translation mistakes can propagate, producing systematic failures despite correct semantic structuring.
Planning Limitations. Our framework does not strengthen the base model’s inherent planning ability. When reasoning paths are poorly identified, semantic analysis alone cannot recover performance, particularly on long-horizon datasets such as ProverQA. Moreover, the planner may inadvertently lengthen reasoning trajectories: ProverQA targets 6–9 step solutions, yet our full framework often extends beyond 10 steps, yielding an accuracy of 68.6%. In contrast, removing planning (woPlan) keeps reasoning within the expected range (6–9 steps) and achieves higher accuracy of 75% (Table 3). This suggests that reasoning length has a critical threshold, beyond which accuracy deteriorates.
Verification Inconsistencies. Although occurring at extremely low frequencies, our method occasionally exhibits hallucination during the verification phase. In these instances, despite making correct intermediate judgments, the system produces final verdicts that contradict its own reasoning steps, indicating a contradiction between reasoning processes and output generation.
F.1 Computational Efficiency Analysis
We analyze the computational efficiency of LogicAgent from both configuration-level and stage-level perspectives, focusing on processing time and token consumption across core components.
Time-Accuracy Trade-offs. As shown in Figure 5(b), LogicAgent exhibits a clear trade-off between reasoning accuracy and computational efficiency. Incorporating planning increases per-sample processing time by approximately 78%, but this additional cost yields structured, goal-oriented reasoning trajectories that are essential for tackling complex multi-hop logic tasks. Similarly, FOL translation contributes significantly to reasoning accuracy by enabling symbolic deduction and FOL consistency checking, yet it introduces substantial latency due to its multi-step transformation and verification pipeline. In contrast, relying solely on natural language input offers faster reasoning with reduced cost but sacrifices FOL rigor and precision in logical reasoning. These results underscore the value of structured symbolic reasoning, particularly when accuracy and interpretability are prioritized.
Table 7: Token Consumption Analysis for RepublicQA
Config Token Type Mean Median
Full Prompt 14,402.44 13,866.00
Completion 3,988.79 3,904.50
Total 18,391.23 18,262.50
woFOL Prompt 11,274.76 10,586.00
Completion 2,788.24 2,732.00
Total 14,063.00 13,405.50
woStatement Prompt 13,113.88 12,125.00
Completion 3,472.03 3,309.50
Total 16,585.92 15,606.00
Token Consumption Patterns. We report token consumption patterns for RepublicQA under three configurations (Table 7). On average, the full setting requires 18.4k tokens, while removing the FOL module (woFOL) reduces usage by 23.5% to 14.1k tokens, with prompt and completion tokens decreasing by 21.7% and 30.1%, respectively. Prompt tokens consistently dominate (75–80% of total), reflecting the heavy contextual demands of multi-hop reasoning. The woStatement setting shows the largest variability, indicating that semantic structuring requirements fluctuate substantially across different philosophical queries.
Stage-Level Timing Breakdown.
To understand intra-system efficiency, we analyze stage-wise processing time for LogicAgent’s full configuration on RepublicQA (Table 8). The Logical Reasoning Stage is the dominant computational bottleneck, accounting for 75.1% of total runtime (139.66s). This is attributed to three interacting factors: (1) multi-path execution across $S_{1}$ and $\lnot S_{1}$, effectively doubling reasoning steps; (2) context-to-FOL translation involving semantic disambiguation and quantifier binding; and (3) execution of detailed plans that enforce step-wise logical progression. The Semantic Structuring Stage, responsible for constructing Greimas’ semiotic square, is comparatively efficient (13.6%), as it involves short-form outputs and deterministic linguistic transformations. The Reflective Verification Stage, while occasionally expensive, benefits from its adaptive design. In most cases, it executes lightweight verification; only uncertain or conflicting cases invoke full re-analysis, keeping its average cost low (11.3%).
Implications. LogicAgent’s modular architecture enables flexible control over computational cost. High-cost components such as planning or FOL translation can be disabled in resource-constrained settings, or selectively activated when tasks demand deeper reasoning.
Table 8: Stage-wise processing time analysis for LogicAgent on RepublicQA
Stage Processing Time (s) Percentage (%)
Semantic Structuring 25.31 13.6
Logical Reasoning 139.66 75.1
Reflective Verification 21.08 11.3
Total 186.05 100.0
Appendix G Full Prompting
Below are detailed prompts used by each component in our reasoning framework, including Semantic Structuring, Translator, Planner, Solver, and Reflective Verification. Each prompt is designed to elicit structured outputs from the language model to support the overall reasoning pipeline.
Case 1 Context: Some individuals who possess a particular skill may also be able to perform the opposite action of that skill. ······ Question: Is the statement ’The just man turns out to be a thief’ correct? Constructing Semiotic Square S1 (Assertion): ”The just man turns out to be a thief”
FOL Translation: $\exists x(Just(x)\land Thief(x))$
S2 (Contrary): ”The just man turns out to be a guardian”
FOL Translation: $\forall x(Just(x)\rightarrow\neg Guardian(x))$
[…]
FOL Translation: $\forall x(Just(x)\rightarrow\neg Thief(x))$
[…]
FOL Translation: $\exists x(Just(x)\land Guardian(x))$
Verifying FOL with CFG
[A2] Validation Status: PASSED
- All quantifiers properly scoped
- Predicate structures syntactically correct
- Logical connectives appropriately placed
- Variable bindings consistent
Verifying logical structure Logical Structure Verification: GreimasCoreValid: true - Core opposition identified GreimasCompleteValid: true - Complete square constructed Translator The system processes contextual information and formalizes eleven key premises: 1. Premise 1: “Some individuals skilled at striking may also be able to defend effectively”
FOL: $\exists y\,(SkillfulAt(Strike,y)\land CanPerform(Defend,y))$
2. Premises…
Multi-Perspective Planning, Reasoning Execution, and Verification
Perspective A ($S_{1}$):
The system generates a 7-step reasoning plan and executes it.
Verdict: $S_{1}$ reasoning concludes Uncertain.
Perspective B ($\lnot S_{1}$):
Parallel reasoning is performed for the contradictory proposition.
Verdict: $\lnot S_{1}$ reasoning concludes True.
Direct Resolution:
Inconsistency Detected: $S_{1}=\text{Uncertain}$ and $\lnot S_{1}=\text{True}$ violate logical consistency. Trigger: System enters Quick Reflection.
Quick Reflection Analysis:
S1 Reasoning: Incorrect. $\lnot S_{1}$ Reasoning: Correct.
Reflection Classification:
[…]
Semantic Structuring You are a reasoning expert. Your task is to analyze a logical proposition using the Greimas’ Semiotic Square framework, which decomposes a proposition into four positions: S1 (original statement), S2 (semantic contrary), ¬S1 (negation of S1), and ¬S2 (negation of S2). Core Steps: 1. Extract Core Proposition: If the question asks ”Is the statement ’X’ correct?”, extract $X$ as $S_{1}$. Preserve original wording exactly. 2. Identify Semantic Contrary: Define $S_{2}$ as a proposition that cannot be true simultaneously with $S_{1}$, though both may be false. Priority opposition types include: Moral: just vs. unjust, good vs. evil Behavioral: help vs. harm, benefit vs. hurt Authority: obedience vs. independent judgment 3. Build Semiotic Square: $S_{1}$: Original target proposition $S_{2}$: Semantic contrary to $S_{1}$ $\lnot S_{1}$: Logical negation of $S_{1}$ $\lnot S_{2}$: Logical negation of $S_{2}$   Example Analysis: Question: Is the statement “repaying a debt is always just” correct? Concept A: just Concept B: unjust $S_{1}$: Repayment of debt is always just.
FOL: $\forall x\,(Debt(x)\land Repaid(x)\rightarrow Just(x))$
$S_{2}$: Repayment of debt is always unjust.
FOL: $\forall x\,(Debt(x)\land Repaid(x)\rightarrow Unjust(x))$
…… $S_{2}$ Type: Contrary
  Output Format (JSON):
  "concept\_A": "...",
  "concept\_B": "...",
  "S1": $\{$"statement": "...", "FOL": "..."\},
  "S2": $\{$"statement": "...", "FOL": "..."\},
  "not\_S1": $\{$"statement": "...", "FOL": "..."\},
  "not\_S2": $\{$"statement": "...", "FOL": "..."\},
Now analyze the following statement using this framework. Question: {question}
Translator You are a logical reasoning expert skilled in translating natural language into precise logical structure. Your task is to extract a list of key premises from the following context. Each premise must be expressed in two formats: 1. A concise and accurate natural-language statement 2. Its corresponding First-Order Logic (FOL) expression written in standard predicate logic FOL rules: Logical conjunction of $expr_{1}$ and $expr_{2}$:  $expr_{1}\land expr_{2}$ Logical disjunction of $expr_{1}$ and $expr_{2}$:  $expr_{1}\lor expr_{2}$ Logical exclusive disjunction of $expr_{1}$ and $expr_{2}$:  $expr_{1}\oplus expr_{2}$ Logical negation of $expr_{1}$:  $\neg expr_{1}$ $expr_{1}$ implies $expr_{2}$:  $expr_{1}\rightarrow expr_{2}$ $expr_{1}$ if and only if $expr_{2}$:  $expr_{1}\leftrightarrow expr_{2}$ Logical universal quantification:  $\forall x$ Logical existential quantification:  $\exists x$   Conventions & Guidelines Use explicit action variables (a) for actions like “repaying” or “obeying”, and object variables (x) for debts, obligations, or rules. Use person or role variables (y) for entities like people, rulers, citizens, friends. Predicates must apply directly to valid entities or actions — never nest predicates: Typed variables:
x $\rightarrow$ debt / obligation / rule
a $\rightarrow$ action
y $\rightarrow$ person / social role (e.g., friend, ruler, citizen)
Focus on extracting premises related to obligation, justice, causality, moral norms. Quantifiers:
$\forall$ (for all), $\exists$ (there exists), and treat Most / Typically as $\forall$ (general statements).
If the context suggests a causal chain (e.g., problematic debt $\rightarrow$ harm $\rightarrow$ unjust), write each causal link as a separate premise — do not collapse into a single line.
  Below is the information you need to deal with right now. Context: {context}  Return your answer in exactly this JSON format:
  "premises": [
      "statement": "...",
      "FOL": "..."
Planner You are a logical reasoning expert. Your task is to draft a step-by-step reasoning plan to determine whether a given logical statement is true, false, or uncertain. The definition of the three options are: True: If the premises can infer the question statement under FOL reasoning rule False: If the premises can infer the negation of the question statement under the FOL reasoning rule Uncertain: If the premises cannot infer whether the question statement is true or false. What to do: 1. Identify the goal (the statement to evaluate). 2. Identify which premises, rules, or definitions are relevant. 3. Break down how to logically connect premises to reach intermediate reasonings. 4. Organize the reasoning steps clearly and sequentially. 5. End with a final step: determine whether the statement in the goal is true or false or uncertain, without making the judgment.   Below is an example Question:
“Repaying one’s debts is always just.”,
$\forall x$ (Debt(x) $\land$ Repaid(x) $\rightarrow$ Just(x))” Premises: Justice involves doing good to friends. FOL: $\forall a$ (Just(a) $\rightarrow$ $\forall y$ (Friend(y) $\rightarrow$ Beneficial(a,y))) ……
  "plan": [
    "Step 1: Identify the goal...
    "Step n: Search for counterexamples...
    "Final Step: Decide whether the premises ...
  Below are the premises and questions you need to derive a plan to solve, please follow the instruction and example aforementioned. Input: Question
{target_statement}
Premises:
{premises}
  Plan: Make sure you only derive the plan. Do not solve the question and do not determine the truth value of the conclusion at the planning stage. This plan will be used to help guiding a language model to follow step-by-step. The expected final step in the plan is to determine whether the the conclusion is true/false/uncertain. Do not solve the question and do not determine the truth value at this stage. Only generate a detailed reasoning plan.
Solver The task is to determine whether the value of the conclusion/question is true/false/uncertain based on the premises. You must refer to the following first-order logic reasoning rules when making logical reasoning. Input Information: 1. Semiotic Square (The statement you need to reason to judge) 2. Formal Premises extracted from the context Your goal is to evaluate whether the statement in the goal logically follows from the premises. Analyze step-by-step. Please solve the question step by step. During each step, please indicate what first-order logic reasoning rules you used. Besides, show the reasoning process by the logical operators including but not limited to: $\oplus$ (either or), $\vee$ (disjunction), $\wedge$ (conjunction), $\rightarrow$ (implication), $\forall$ (universal), $\exists$ (existential), $\neg$ (negation), $\leftrightarrow$ (equivalence). You can combine natural language and logical operators when doing reasoning.  Definitions: True: A statement is “true” if it necessarily follows from the given premises using logical rules. False: A statement is “false” if it is contradicted by the premises or its negation is logically inferred from them or if there are counterexamples. Uncertain: A statement is “uncertain” if there is insufficient information in the premises to determine its truth value conclusively.   Now analyze input Goal: {target_statement} Premises: {premises} Plan: {PLAN}  Output JSON Format (place this at the end, Ensure the JSON is valid (no trailing commas)):
  "steps": [
    "Step 1: ...",
    "Step 2: ...",
    "Final answer: {true/false/uncertain}"
  "verdict": "True" | "False" | "Uncertain"
[…]
  "verdict": "True" | "False" | "Uncertain",
  "reason": "Type 1: S1 reasoning correct → Return S1’s verdict"|
  "Type 2: S1 incorrect, ¬S1 correct with Uncertain verdict → Return Uncertain"|
  "Type 3: S1 correct with Uncertain verdict → Return Uncertain" |
  "Type 4: S1 incorrect, ¬S1 correct with True verdict → Return False" |
  "Type 5: S1 incorrect, ¬S1 correct with False verdict → Return True" |
  "Type 6: Both S1 and ¬S1 incorrect → Return independently verified result"
  Verification Execution: Original Execution: [[EXECUTION]] Verify: Please indicate the revised answer at the end using CURLY BRACKETS. The response must be one of:
  "verdict": "True" | "False" | "Uncertain",
  "reason": "Type 1: S1 reasoning correct → Return S1’s verdict"|
  "Type 2: S1 incorrect, ¬S1 correct with Uncertain verdict → Return Uncertain"|
  "Type 3: S1 correct with Uncertain verdict → Return Uncertain" |
  "Type 4: S1 incorrect, ¬S1 correct with True verdict → Return False" |
  "Type 5: S1 incorrect, ¬S1 correct with False verdict → Return True" |
  "Type 6: Both S1 and ¬S1 incorrect → Return independently verified result"